123456789234:

a car Travels 368 miles in five hours and 45 minutes. How many miles does it travel per hour

Extrinix:

What do you think the answer is?

Extrinix:

@123456789234

Extrinix:

First off, find each of the values, then, convert the time into one time (\(minutes\) is preferred) divide \(368\) by the amount of time in minutes, take your answer and multiply it by \(60\) you'll have your \(answer\).

jhonyy9:

so like a first step we need convert the 5 hours in minutes for what we can use the rule of 3 simple and write in this way 1 hour mean 60 minutes 5 hours mean x cross multiplie and calcule the x value x = 5*60/1 = 5*60 = 300 min. using this above wrote we know that this car travels 368 miles in 345 minutes we need to find how many miles does it travel per hour to find this we can use the rule of 3 simple and write in this way 368 miles travels in 345 minutes x miles travels in 60 minutes what equal 1 hour - cross multiplie and calcule the x value x = 368*60/345 = ? miles/hour hope helped understandably easy @AZ opinon pls - ty

Extrinix:

@jhonyy9 Thanks!

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Extrinix @jhonyy9 Thanks! \(\color{#0cbb34}{\text{End of Quote}}\) np yw